OpenStudy (anonymous):

SOLVE AND PLEASE CHECK.LABEL ANY CONTRADICTIONS OR IDENTITIES. 5(3-3T)=30 DONT FORGET TO PROVIDE THE CHECK THE PROBLEM

jimthompson5910 (jim_thompson5910):

5(3-3T)=30 15-15T=30 -15T=30-15 -15T=15 T=15/(-15) T=-1 ------------------------------------------------------- Check 5(3-3T)=30 5(3-3(-1))=30 5(3+3)=30 5(6)=30 30=30 So the solution T = -1 is verified

OpenStudy (anonymous):

THANKS I DID IT RIGHT I KEEP FORGET TO PUT MY SIGNS IN THE ANSWER